A spherical raindrop (r = 0.0015 m) falls from a cloud. The drag coefficient is 0.60. The density of the water is 1000 kg/m^3 and the density of the air it falls through is 1.2 kg/m^3. The shape of the drop doesn’t change during the fall, and the terminal velocity is 7.3 m/s. At this terminal speed, what is the magnitude of the resistive force acting on the drop?  
Answer #1

magnitude of resistive force

F = 0.5 * rho * A * C v^2

F = 0.5* 1.2* (0.0015)^2* 0.6*7.3^2

F =1.355* 10^-4 N

=======

Comment before rate in case any doubt, will reply for sure.. goodluck